# Simple Yet Effective Selective Imputation for Incomplete Multi-view Clustering

### v1

@@ Abstract
Incomplete multi-view data, where different views suffer from missing and unbalanced observations, pose significant challenges for clustering. Existing imputation-based methods attempt to estimate missing views to restore data associations, but indiscriminate imputation often introduces noise and bias, especially when the available information is insufficient. Imputation-free methods avoid this risk by relying solely on observed data, but struggle under severe incompleteness due to the lack of cross-view complementarity. To address this issue, we propose Informativeness-based Selective imputation Multi-View Clustering (ISMVC). Our method evaluates the imputation-relevant informativeness of each missing position based on intra-view similarity and cross-view consistency, and selectively imputes only when sufficient support is available. Furthermore, we integrate this selection with a variational autoencoder equipped with a mixture-of-Gaussians prior to learn clustering-friendly latent representations. By performing distribution-level imputation, ISMVC not only stabilizes the aggregation of posterior distributions but also explicitly models imputation uncertainty, enabling robust fusion and preventing overconfident reconstructions. Compared with existing cautious imputation strategies that depend on training dynamics or model feedback, our method is lightweight, data-driven, and model-agnostic. It can be readily integrated into existing IMC models as a plug-in module. Extensive experiments on multiple benchmark datasets under a more realistic and challenging unbalanced missing scenario demonstrate that our method outperforms both imputation-based and imputation-free approaches. The code and datasets are available at: https://github.com/Starnever0/ISMVC.
@@ I Introduction
Multi-view data has attracted increasing attention for its ability to provide more comprehensive descriptions of real-world entities or events. For example, on social media platforms, the user generated content may be composed of text view (e.g., descriptions or comments), image view (e.g., uploaded photos), and location view (e.g., GPS-based check-ins). These views collectively depict user behaviors: text reveals opinions or sentiments, images convey visual content, and location provides spatial context. By leveraging the consistency and complementarity across views, multi-view learning has achieved remarkable success in downstream tasks such as classification [hanTrusted2021, seeland2021multi], clustering [luoSimple2024, moujahid2025advanced], multimodal large language models [yin2024survey] and embodied intelligence [cangelosi2015embodied]. However, due to various practical limitations, real-world multi-view data are often incomplete, causing most existing methods to inevitably degrade or even fail. In this paper, we focus on the Incomplete Multi-view Clustering (IMC) problem, which integrates incomplete views to help identify essential grouping structure in an unsupervised manner.
Figure 1: Motivation. Clustering accuracy peaks when partially imputing well-supported positions, while full or no imputation leads to suboptimal results.
@@
A predominant line of research follows an imputation-based paradigm for the IMC problem where the missing views are estimated based on the available ones to enhance data association [wang2018partial, li2014partial]. While such approaches improve view completeness, their effectiveness highly depends on the reliability of the imputed information. Indiscriminate imputation of all missing positions, without assessing the reliability of reference data, can inject unreliable pseudo-data and distort latent representations, especially under high missing rates or unbalanced view conditions [yuanPrototype2025, fang2021unbalanced]. For instance, in user-generated content, the location view frequently exhibit higher missing rates due to privacy settings or device limitations, resulting in information scarcity when intra-class neighbors are also absent. Blindly filling such positions introduces noisy information that destabilizes latent inference, disrupts structural consistency, and ultimately degrades clustering performance.
Different from the aforementioned imputation-based IMC methods, another line of imputation-free methods aim to directly learn clustering representations from available views, avoiding the potential bias introduced by inaccurate imputation [xu2022deep, baiGraphguided2024]. However, these methods, such as view-weighted fusion or shared latent space modeling, also struggle addressing highly unbalanced or severely incomplete scenarios. The lack of complementary information from missing views weakens cross-view correspondence, leading to representational collapse and degraded clustering performance. Fig. 1 demonstrates this phenomenon: while no imputation (0%) leads to information loss, excessive imputation (100%) introduces biases and degrades performance. The optimal performance is achieved through selective imputation, validating our motivation for informativeness-based imputation selection strategies.
To mitigate the adverse effects of inaccurate imputations, various strategies have been proposed to promote more cautious and reliable use of imputed data. For example, some methods introduce adaptive mechanisms which assessing imputation quality via loss changes or clustering consistency, and accordingly reduce the influence of potentially misleading samples by down-weighting or excluding them [tangDeepSafeIncomplete2022, puAdaptive2024b]. While these methods improve robustness to some extent, they still suffer from critical limitations. First, relying on the training dynamics of specific models to assess imputation quality introduces significant variability and unreliability in decision-making. When models are inadequately trained, inherently unstable, or produce erroneous clustering outcomes, the resulting quality evaluations are compromised, leading to misguided or underutilized imputations. Second, most of them rely on intricate dynamic procedures—such as bi-level optimization, which incurs substantial computational overhead. This severely restricts scalability and poses significant barriers to efficient deployment in large-scale or resource-limited environments.
To address the above limitations, we propose a lightweight, data-driven selective imputation framework for the IMC problem. Rather than blindly imputing all missing positions, our method selectively imputes only well-supported positions by quantifying their imputation-relevant informativeness from intra-view and cross-view sources. Furthermore, to mitigate the bias introduced by unreliable imputations, we incorporate a distribution-level strategy within a variational inference framework, where imputation is modeled with explicit variance reflecting uncertainty. This uncertainty-aware design allows the product-of-experts fusion to automatically down-weight unreliable views, thereby enhancing robustness and semantic alignment. Extensive experiments demonstrate the superiority of our method over existing approaches and validate the generality of the proposed Informativeness-Based Selective Imputation (IBSI) strategy across various clustering baselines. Our main contributions are summarized as follows:
  • From the perspective of imputation decision-making, we propose a simple yet effective selective imputation strategy to address the trade-off between imputation utility and imputation risk. To the best of our knowledge, this is the first attempt in the IMC literature to perform imputation selectively at the level of individual missing positions.
  • We introduce a prior assessment of imputation uncertainty that operates at the data level, independent of specific models or clustering tasks. This approach enhances robustness and adaptability across diverse datasets and scenarios, while maintaining computational efficiency for practical deployment.
@@
  • We design a distribution-level imputation strategy within a variational inference framework that exploits neighborhood-based sample variance for imputation uncertainty quantification, implicitly down-weighting low-quality imputations during fusion to mitigate biases and enhance the robustness of shared representations.
@@ II Related Works
To tackle the challenges posed by incomplete multi-view data, existing IMC methods generally follow two main directions: (1) Imputation-based methods. Classical approaches leverage statistical correlations within and across views to estimate missing data [tran2017missing, liuEfficient2019], or explicitly learning view-translation functions to recover the missing view from available ones [linCOMPLETERIncompleteMultiView2021] . With the rise of deep generative models, a variety of methods have employed generative adversarial networks (GANs) [shang2017vigan, wang2018partial, zhang2020deep], variational autoencoders (VAEs) [xuAudioVisual2024], and more recently, diffusion models [wenDiffusionbased2024] to synthesize missing data. (2) Imputation-free methods. In contrast, Imputation-free methods focus on designing missing-aware multi-view fusion or alignment strategies to directly learn the clustering structure from only the observed parts. For example,  [tengURRLIMVC2024] introduces a unified representation learning framework based on an attention-driven autoencoder. [xuAdaptive2023] avoids view imputation by learning a shared latent space through adaptive feature projection, which dynamically adjusts projection weights based on view quality and availability. While both paradigms have shown promise, they often suffer from performance degradation under severe or unbalanced missing conditions, as discussed in the Introduction. Our work departs from these extremes by proposing a lightweight, informativeness-based strategy that selectively imputes only reliable missing positions.
@@
Except of exploring more accurate impute method, several imputation-based studies adopt more cautious training schemes to alleviate negative effects from inaccurate imputation. One representative direction is the use of complete-case analysis: models are trained only on fully observed samples, while missing views are estimated or evaluated through auxiliary prediction networks [linDual2022]. However, this method shares the limitations of imputation-free approaches, often yielding suboptimal performance in extreme scenarios. Another line of research employs more sophisticated techniques, involving imputation uncertainty evaluation during training and downstream performance consistency after imputing [tangDeepSafeIncomplete2022, yanDeep2025]. For instances, [xieExploring2023] model each missing data with a distribution conditioning on the available views and thus introducing uncertainty, evidence-based fusion strategy are employed to adaptively utilize the uncertainty. [huangUnified2024a] use the half-quadratic minimization technique automatically weight different instances, alleviating the impact of outliers and unreliable restored data. More recently,  [hu2024reliable] introduces a reliable imputation guidance module that distinguishes intrinsic zeros from technical zeros using cluster-level confidence (based on zero rates, means, and variances), selectively imputing only the latter while jointly reconstructing instance- and feature-level structures.  [xi2025lrgr] proposes a two-stage quality-adaptive framework: a local refinement stage using cross-view contrastive learning and view-specific prototypes, followed by a global realignment stage with confidence-weighted pseudo-labels to mitigate distribution shifts from unreliable imputation. For downstream result, [puAdaptive2024b] perform imputation followed by an evaluation of clustering consistency between imputed and original data, discarding imputed data if clustering performance degrades and reverting to training on the original, non-imputed data. However, these methods either rely on model-dependent training dynamics for quality assessment or impose complex architectural constraints, resulting in limited granularity and scalability. This motivates us to propose a more general, lightweight, and fine-grained framework that explicitly quantifies position-level information availability to safely leverage imputed data.
@@ II-C Variational Autoencoder
Variational Autoencoders (VAEs) [kingma2013auto] have emerged as a generative framework with explicit uncertainty modeling, and they have been widely adopted in representation learning and clustering. By adopting diverse priors over latent representations and incorporating different generative assumptions, classical works have enhanced the clustering ability and discriminative power of VAEs, for example through Mixture-of-Gaussians (MoG) prior assumption [jiangVariational2017a, dilokthanakul2016deep], graph-based constraints [yang2019deep], or disentangled continuous and discrete factors [dupont2018learning].
In multi-view settings, the probabilistic nature of VAEs makes them particularly suitable for encoding different modalities into view-specific latent spaces and fusing them through reasonable principles. For example, suzuki et al. [suzukiJoint2016] introduced a joint multimodal VAE that enforces a shared latent representation across views, enabling coherent generation and cross-modal inference. He et al. [heM^2VAE2025] leverages type-specific latent variables and a Product-of-Experts (PoE) fusion to disentangle common and modality-specific representations. VariGANs [zhaoMultiView2018] combine variational inference with GANs, where VAEs capture cross-view global appearance and GANs refine details for high-quality multi-view image generation. Yin et al. [yinShared2020] introduced adaptive weights together with a MoG prior to encourage clustering-friendly shared representations. Another important direction partitions the latent space into shared and private components, effectively disentangling cross-view commonality from view-specific complementary information [xu2021multi].
For real-world scenarios with partially observed views, VAEs have been extended to incomplete multi-view learning and imputation. MultiVI [ashuach2023multivi] integrates partially missing multi-omics data by learning a unified latent space, enabling joint analysis despite incomplete modalities. CMVAE [caiRealize2024b] introduces a MoG latent representation that jointly generates all views, enabling posterior-based recovery of missing views and improving clustering and classification performance. Xu et al. [xu2024deep] address incomplete multi-view clustering by employing view-specific VAEs and aggregating their latent representations via a Product-of-Experts to obtain a robust shared representation.
@@
Figure 2: Overview of the ISMVC framework. An informativeness-based imputation position selection mechanism scores missing positions $\tilde{\bm{x}}_{i}^{v}$ based on intra-view similarity and cross-view correlation, imputing only those with sufficient support. Variational inference uses Product-of-Experts to aggregate view posteriors, with distribution-level imputation of missing views via latent neighbors. In the generative process, each observed view is generated from the cluster-aware latent representation $\bm{z}$.
@@ III Method
In this section, we first outline the generative assumptions that form the foundation of our approach. We then introduce the Informativeness-Based Imputation Position Selection mechanism, which assigns an information score to each missing position to determine whether imputation is feasible. Next, we detail the imputation strategy in a variational inference framework. The overall architecture of our Informativeness-based Selective imputation Multi-View Clustering (ISMVC) method is illustrated in Fig. 2.
@@ III-A Notations and Problem Statement
Let $\{\{\bm{x}_{i}^{v}\}_{v=1}^{V}\}_{i=1}^{N}$ represent an incomplete multi-view dataset consisting of $N$ samples and $V$ views. Each $\bm{x}_{i}^{v}\in\mathbb{R}^{d_{v}}$ denotes the feature of the $i$-th sample in view $v$, with $d_{v}$ being its feature dimension. The ground-truth label of sample $i$ is denoted by $y_{i}\in\{1,\dots,K\}$, where $K$ is the number of classes. We define a binary indicator matrix $\mathbf{M}\in\{0,1\}^{N\times V}$ to describe the observed pattern, where $\mathbf{M}_{i,v}=1$ if $\bm{x}_{i}^{v}$ is observed, and $\mathbf{M}_{i,v}=0$ otherwise. For notational convenience, we denote a missing position as $\tilde{\bm{x}}_{i}^{v}$ if $\mathbf{M}_{i,v}=0$. The set of observed views for sample $i$ is defined as $\mathcal{V}_{i}=\{v\mid\mathbf{M}_{i,v}=1\}$ and the set of observed samples for view $v$ is defined as $\mathcal{X}^{v}=\{\bm{x}_{i}^{v}\mid\mathbf{M}_{i,v}=1\}$. The goal of IMC is to cluster the $N$ samples into $K$ groups based on incomplete multi-view representations.
III-B Deep Multi-view Gaussian Mixture Model
To leverage the clustering capability of Gaussian Mixture Models (GMMs), we incorporate a GMM as the prior in the Variational Autoencoder (VAE). Following the Deep Multi-view Gaussian Mixture Model (DMGMM) assumption in [jiangVariational2017a, xu2024deep], the generative process proceeds in three steps: First, a shared discrete cluster assignment $\bm{c}\in\{0,1\}^{K}$ is drawn from a categorical distribution with prior probabilities $\bm{\pi}\in\mathbb{R}^{K}$, where each $\pi_{k}$ denotes the prior probability of the $k$-th cluster and satisfies $\sum_{k=1}^{K}\pi_{k}=1$. Then, a continuous latent variable $\bm{z}\in\mathbb{R}^{d}$ is sampled from a Gaussian distribution conditioned on the cluster assignment $\bm{c}$. Finally, given $\bm{z}$, each view-specific observation is independently generated from its corresponding conditional distribution. Under this assumption, the joint probability of an incomplete multi-view sample ($i$-th sample) can be written as:

### v2

@@ Abstract
Incomplete Multi-view Clustering (IMC) has emerged as a significant challenge in multi-view learning. A predominant line for IMC is data imputation; however, indiscriminate imputation can result in unreliable content. Recently, researchers have proposed selective imputation methods that use a post-imputation assessment strategy: (1) impute all or some missing values, and (2) evaluate their quality through clustering tasks. We observe that this strategy incurs substantial computational complexity and is heavily dependent on the performance of the clustering model. To address these challenges, we first introduce the concept of pre-imputation assessment. We propose an Implicit Informativeness-based Selective Imputation (SI3) method for incomplete multi-view clustering, which explicitly addresses the trade-off between imputation utility and imputation risk. SI3 evaluates the imputation-relevant informativeness of each missing position in a training-free manner, and selectively imputes data only when sufficient informative support is available. Under a multi-view generative assumption, SI3 further integrates selective imputation into a variational inference framework, enabling uncertainty-aware imputation at the latent distribution level and robust multi-view fusion. Compared with existing selective imputation strategies, SI3 is lightweight, data-driven, and model-agnostic, and can be seamlessly incorporated into existing incomplete multi-view clustering frameworks as a plug-in strategy. Extensive experiments on multiple benchmark datasets demonstrate that SI3 consistently outperforms both imputation-based and imputation-free methods, particularly under challenging unbalanced missing scenarios. The code and datasets are available at: https://github.com/Starnever0/SI3.
@@ I Introduction
Multi-view data has attracted increasing attention for its ability to provide more comprehensive descriptions of real-world entities or events. For example, on social media platforms, the user generated content may be composed of text view (e.g., descriptions or comments), image view (e.g., uploaded photos), and location view (e.g., GPS-based check-ins). These views collectively depict user behaviors: text reveals opinions or sentiments, images convey visual content, and location provides spatial context. By leveraging the consistency and complementarity across views, multi-view learning has achieved remarkable success in downstream tasks such as classification [11, 28], clustering [22, 18, 24, 25], multimodal large language models [47] and embodied intelligence [5]. However, due to various practical limitations, real-world multi-view data are often incomplete, causing most existing methods to degrade or even fail inevitably. In this paper, we focus on the Incomplete Multi-view Clustering (IMC) problem, which integrates incomplete views to help identify essential grouping structure in an unsupervised manner.
Figure 1: The clustering accuracies of three representative IMC methods enhanced by the proposed pre-imputation assessment strategy, which prioritizes imputing more informative missing positions in a training-free manner.
@@
A predominant line to solve this problem is imputation-based methods [34, 38], which estimate the missing positions via related positions (such as the positions of the same sample or cluster). While these methods enhance view completeness, their effectiveness highly depends on the reliability of the imputed data. However, indiscriminate imputation of all missing positions, without assessing the imputation-relevant informativeness provided by related observations related positions would involve unreliable content, especially under high missing rates [48] or unbalanced view conditions [8]. For instance, in user-generated content, the location view frequently exhibits higher missing rates due to privacy settings or device limitations, resulting in information scarcity when intra-cluster neighbors are also absent. In this condition, imputing such positions will introduce noisy information.
Recently, some researchers recognize this issue and propose selective imputation methods [31, 27] that utilize a post-imputation assessment strategy. These methods first impute all or partial positions, followed by clustering on the new data. They then evaluate the quality of the imputation through changes in loss or clustering consistency and reduce the influence of potentially noisy imputed positions by down-weighting or excluding them. This process is repeated multiple times until the model achieves optimal results. These methods have achieved good results, but this post-imputation assessment strategy still faces two new challenges: (1) the additional post-imputation assessment modules impose substantial computational complexity; (2) the post-imputation assessment strategy relies heavily on the clustering model to evaluate imputation quality, making the assessment sensitive to initialization, model stability, and intermediate training outcomes. When the clustering model is insufficiently trained, unstable, or exhibits poor performance, the imputation assessment becomes unreliable.
In the paper, we propose a simple yet effective selective imputation method (Implicit Informativeness-based Selective Imputation, SI3) for this problem. The key motivations are: (1) not all positions should be imputed; (2) the decision to impute a position depends on the informativeness that can facilitate imputation for that position. Based on this, we introduce a pre-imputation assessment strategy, which uses a training-free manner to quantify the position’s imputation-relevant informativeness from both intra-view and cross-view sources. In addition, to further mitigate the noise caused by unreliable imputations, we propose a distribution-level imputation strategy within a variational inference framework. The variance of the imputed distributions captures the uncertainty of imputations. This uncertainty-aware approach allows the multi-view fusion process to automatically down-weight unreliable views, thereby improving both robustness and semantic alignment across views.
Extensive experiments demonstrate that our method consistently outperforms existing approaches across various missing rates. Fig. 1 shows an intuitive example: we enhance three representative IMC methods by the informativeness-based imputation position selection mechanism in our SI3. The imputation ratio is gradually increasing from 0% to 100%. Experimental results indicate that both under-imputation and over-imputation yield poor outcomes: under-imputation causes information loss, while over-imputation introduces noise and degrades performance. Optimal performance is achieved at a moderate imputation ratio, indicating only partial imputation is beneficial. By identifying and imputing only the missing positions with high imputation-relevant informativeness, SI3 effectively attains this favorable imputation regime and achieves superior performance. Our main contributions are summarized as follows:
  • We have observed that selective imputation contributes significantly to addressing the important IMC problem (Fig. 1). Moreover, we first introduce the concept of pre-imputation assessment to achieve a simple yet efficient selection of imputation positions.
  • The proposed pre-imputation assessment strategy operates at the data level, independent of specific models or clustering tasks. When integrated as a plug-in into representative baselines such as DCP[20], it incurs less than 2% additional runtime overhead while achieving up to 10.48% improvement in clustering accuracy, demonstrating strong robustness, adaptability, and practical efficiency across different baselines.
@@
  • We design a distribution-level imputation strategy within a variational inference framework that models imputation uncertainty, implicitly down-weighting low-quality imputations during fusion to improve robustness.
  • Extensive experiments on diverse benchmark datasets demonstrate the effectiveness and generality of the proposed SI3 framework. In particular, under severe and imbalanced missingness with an overall missing rate of 50%, SI3 achieves a 12.7% improvement in clustering accuracy over the imputation-free baseline DVIMC[39].
@@ II Related Works
To tackle the challenges posed by incomplete multi-view data, existing IMC methods generally follow two main directions: (1) Imputation-based methods. Classical approaches leverage statistical correlations within and across views to estimate missing data [33, 23], or explicitly learning view-translation functions to recover the missing view from available ones [21]. With the rise of deep generative models, a variety of methods have employed generative adversarial networks (GANs) [29, 34, 49], variational autoencoders (VAEs) [43], and more recently, diffusion models [35] to synthesize missing data. (2) imputation-free methods. In contrast, Imputation-free methods focus on designing missing-aware multi-view fusion or alignment strategies to directly learn the clustering structure from only the observed parts [41, 3]. For example,  [32] introduces a unified representation learning framework based on an attention-driven autoencoder. [40] avoids view imputation by learning a shared latent space through adaptive feature projection, which dynamically adjusts projection weights based on view quality and availability.
While both paradigms have shown promise, they often suffer from performance degradation under severe or unbalanced missing conditions, as discussed in the Introduction. Our work departs from these extremes by proposing a lightweight, informativeness-based strategy that selectively imputes only reliable missing positions.
@@
Except of exploring more accurate impute method, several studies adopt more cautious training schemes to alleviate negative effects from inaccurate imputation. One representative direction is the use of complete-case training paradigm: models are trained only on fully observed samples, while missing views are estimated at evaluation time using auxiliary prediction networks [20]. However, this method shares the limitations of imputation-free approaches, often yielding suboptimal performance in extreme scenarios. Another line of research employs more sophisticated techniques, involving explicitly evaluate imputation reliability after imputation by monitoring training dynamics and downstream performance [44]. For instances,  [31] propose a bi-level optimization framework in which missing views are dynamically imputed from semantic neighbors, and the quality of these imputations is assessed through variations in the training loss. Based on this assessment, imputed samples that are deemed unreliable are adaptively down-weighted or excluded from subsequent optimization, thereby reducing the risk of performance degradation caused by inaccurate imputations. Similarly, [14] use the half-quadratic minimization technique automatically weight different samples, alleviating the impact of outliers and unreliable restored data. More recently,  [13] introduces a reliable imputation guidance module that distinguishes intrinsic zeros from technical zeros using cluster-level confidence (based on zero rates, means, and variances), selectively imputing only the latter while jointly reconstructing sample- and feature-level structures.  [36] proposes a two-stage quality-adaptive framework: a local refinement stage using cross-view contrastive learning and view-specific prototypes, followed by a global realignment stage with confidence-weighted pseudo-labels to mitigate distribution shifts from unreliable imputation. For downstream result,  [27] adopt a cluster-oriented latent-space imputation strategy, followed by a post-imputation evaluation of clustering outcomes. When the representations obtained after imputation yield inferior clustering results compared to those learned from unimputed data, the imputation process is discarded and subsequent training relies solely on the original representations.
Despite their effectiveness, these methods predominantly adopt post-imputation assessment paradigms, where imputation quality is assessed only after missing positions have already been filled and incorporated into training. Such strategies either rely heavily on model-dependent training dynamics or introduce complex architectural constraints, resulting in limited granularity, scalability, and robustness. These limitations motivate us to propose a more general and lightweight pre-imputation assessment framework, which explicitly quantifies position-level imputation-relevant informativeness at the data level and selectively determines whether a missing position should be imputed before training, enabling safer and more principled utilization of imputed data.
@@ II-C Variational Autoencoder
Variational Autoencoders (VAEs) [17] have emerged as a generative framework with explicit uncertainty modeling, and they have been widely adopted in representation learning and clustering. By adopting diverse priors over latent representations and incorporating different generative assumptions, classical works have enhanced the clustering ability and discriminative power of VAEs, for example through Mixture-of-Gaussians (MoG) prior assumption [15, 6], graph-based constraints [45], or disentangled continuous and discrete factors [7].
In multi-view settings, the probabilistic nature of VAEs makes them particularly suitable for encoding different modalities into view-specific latent spaces and fusing them through reasonable principles. For example, suzuki et al. [30] introduced a joint multimodal VAE that enforces a shared latent representation across views, enabling coherent generation and cross-modal inference. He et al. [12] leverages type-specific latent variables and a Product-of-Experts (PoE) fusion to disentangle common and modality-specific representations. VariGANs [50] combine variational inference with GANs, where VAEs capture cross-view global appearance and GANs refine details for high-quality multi-view image generation. Yin et al. [46] introduced adaptive weights together with a MoG prior to encourage clustering-friendly shared representations. Another important direction partitions the latent space into shared and private components, effectively disentangling cross-view commonality from view-specific complementary information [42].
For real-world scenarios with partially observed views, VAEs have been extended to incomplete multi-view learning and imputation. MultiVI [2] integrates partially missing multi-omics data by learning a unified latent space, enabling joint analysis despite incomplete modalities. CMVAE [4] introduces a MoG latent representation that jointly generates all views, enabling posterior-based recovery of missing views and improving clustering and classification performance. Xu et al. [39] address incomplete multi-view clustering by employing view-specific VAEs and aggregating their latent representations via a Product-of-Experts to obtain a robust shared representation.
@@
Figure 2: Overview of the SI3 framework. An informativeness-based imputation position selection mechanism scores missing positions $\tilde{\bm{x}}_{i}^{v}$ based on intra-view similarity and cross-view correlation, imputing only those with sufficient support. Variational inference uses Product-of-Experts to aggregate view posteriors, with distribution-level imputation of missing views via latent neighbors. In the generative process, each observed view is generated from the cluster-aware latent representation $\bm{z}$.
@@ III Method
In this section, we present the Implicit Informativeness-based Selective Imputation (SI3) method in detail, together with its implementation.
@@ III-A Notations and Problem Statement
Let $\{\{\bm{x}_{i}^{v}\}_{v=1}^{V}\}_{i=1}^{N}$ represent an incomplete multi-view dataset consisting of $N$ samples and $V$ views. Each $\bm{x}_{i}^{v}\in\mathbb{R}^{d_{v}}$ denotes the feature of the $i$-th sample in view $v$. $d_{v}$ is the feature dimension. The ground-truth label of the $i$-th sample is denoted by $y_{i}\in\{1,\dots,K\}$, where $K$ is the number of classes. We define a binary indicator matrix $\mathbf{M}\in\{0,1\}^{N\times V}$ to describe the observed pattern, where $\mathbf{M}_{i,v}=1$ if $\bm{x}_{i}^{v}$ is observed, and $\mathbf{M}_{i,v}=0$ otherwise. For notational convenience, we denote a missing position as $\tilde{\bm{x}}_{i}^{v}$ if $\mathbf{M}_{i,v}=0$. The set of observed views for the $i$-th sample is defined as $\mathcal{V}_{i}=\{v\mid\mathbf{M}_{i,v}=1\}$ and the set of observed samples for view $v$ is defined as $\mathcal{X}^{v}=\{\bm{x}_{i}^{v}\mid\mathbf{M}_{i,v}=1\}$. The goal of IMC is to cluster the $N$ incomplete samples into $K$ groups.
III-B The SI3 Framework
In this section, we first outline the generative assumptions that form the foundation of the SI3. We then introduce the pre-imputation assessment strategy, which calculates an informativeness score to each missing position to determine whether imputation is feasible. Finally, we detail the imputation strategy in a variational inference framework. The overall architecture of the SI3 is illustrated in Fig. 2.
III-B1 The Multi-view Generative Assumption
In this paper, we introduce a multi-view generative assumption, which follows the formulation of the Deep Multi-view Gaussian Mixture Model (DMGMM) and consist of a three-step generative process: first, a shared discrete cluster assignment $\bm{c}\in\{0,1\}^{K}$ is drawn from a categorical distribution with prior probabilities $\bm{\pi}\in\mathbb{R}^{K}$, where each $\pi_{k}$ denotes the prior probability of the $k$-th cluster and satisfies $\sum_{k=1}^{K}\pi_{k}=1$. Then, a continuous latent variable $\bm{z}\in\mathbb{R}^{d}$ is sampled from a Gaussian distribution conditioned on the cluster assignment $\bm{c}$. Finally, given $\bm{z}$, each view-specific observation is independently generated from its corresponding conditional distribution. Under this assumption, the joint probability of an incomplete multi-view sample can be written as: